# ADHint: Adaptive Hints with Difficulty Priors for Reinforcement Learning

### v1

@@ Abstract
To combine the advantages of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), recent methods have integrated “hints” into post-training, which are prefix segments of complete reasoning trajectories, aiming for powerful knowledge expansion and reasoning generalization. However, existing hint-based RL methods typically ignore difficulty when scheduling hint ratios and estimating relative advantages, leading to unstable learning and excessive imitation of off-policy hints. In this work, we propose ADHint, which treats difficulty as a key factor in both hint-ratio schedule and relative-advantage estimation to achieve a better trade-off between exploration and imitation. Specifically, we propose Adaptive Hint with Sample Difficulty Prior, which evaluates each sample’s difficulty under the policy model and accordingly schedules an appropriate hint ratio to guide its rollouts. We also introduce Consistency-based Gradient Modulation and Selective Masking for Hint Preservation to modulate token-level gradients within hints, preventing biased and destructive updates. Additionally, we propose Advantage Estimation with Rollout Difficulty Posterior, which leverages the relative difficulty of rollouts with and without hints to estimate their respective advantages, thereby achieving more balanced updates. Extensive experiments across diverse modalities, model scales, and domains demonstrate that ADHint delivers superior reasoning ability and out-of-distribution generalization, consistently surpassing existing methods in both pass@1 and avg@8. Our code and dataset will be made publicly available upon paper acceptance. *: Equal contribution, 🖂: Corresponding author.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) [yang2024qwen2, yang2025qwen3, li2025system, zeng2025glm] and Multimodal Large Language Models (MLLMs) [bai2025qwen3vltechnicalreport, liu2025visual, huang2025vision, wang2025internvl3] highlight the potential to enhance reasoning capabilities by large-scale Reinforcement Learning with Verifiable Rewards (RLVR) [shao2024deepseekmath, guo2025deepseek, yu2025dapo], such as Proximal Policy Optimization (PPO) [schulman2017proximal] and Group Relative Policy Optimization (GRPO) [shao2024deepseekmath, guo2025deepseek]. However, RLVR primarily reinforces existing behaviors rather than instilling novel reasoning abilities [chusft, li2025system]. On the contrary, Supervised Fine-Tuning (SFT) offers a complementary approach to facilitating the incorporation of new knowledge and reasoning patterns [chen2025sft, yan2025learning]. Nevertheless, the SFT models often exhibit limited generalization to out-of-distribution scenarios. These contrasting characteristics of RL and SFT motivate a promising research direction: How can RL and SFT effectively collaborate to enhance both reasoning abilities and out-of-distribution generalization?
@@
To achieve this goal, early methods [yan2025learning, ma2025learning, dong2025rl, fu2025srft] attempt to incorporate on-policy RL with full reasoning trajectories to internalize external knowledge. However, these methods often struggle with insufficient exploration [wu2025templaterlstructuredtemplateguidedreinforcement, xi2025bapostabilizingoffpolicyreinforcement] and even training collapse [liang2025squeezesoakedspongeefficient, cui2025entropymechanismreinforcementlearning]. Recent methods [Huang_2025_ICCV, liu2025uft, zhang2025stephint] introduce hints, which refer to the prefix segments of complete reasoning trajectories, allowing the policy model to explore the continuation rather than reinforcing the model to mimic the entire sequence. Specifically, these methods assign fixed [zhang2025stephint] or time-varying [liu2025uft, huang2025blending] hint ratios to all samples (or only to the hardest samples [Huang_2025_ICCV, liu2025ghpo]), and then ultize hint-rollouts derived from hint-guided reasoning together with naive-rollouts generated without hints for group relative advantage estimation [zhang2025stephint, huang2025blending].
@@
Despite their empirical successes, existing hint-based RL methods are sensitive to sample-difficulty in the hint-ratio schedule and rollout-difficulty in the relative-advantage estimation. Specifically, most previous methods apply a uniform hint ratio to both easy and hard samples for model reasoning, and then generate hint-rollouts with various difficulties. Policy learning from such heterogeneous rollouts could introduce high-variance model optimization, and thus undermine training stability [bengio2009curriculum, wang2025dump, parashar2025curriculum]. As shown in Figure 1(a), the model with annealing hint ratio (denoted as Baseline) meets with training collapse since its entropy rises sharply at the end of training, indicating insufficient absorption of external knowledge and a drift in the model’s own policy. In contrast, when introducing the hint-ratio schedule with sample-difficulty (our method, discussed later), the entropy converges stably so that the model could learn reasoning ability, not superficial patterns in off-policy hints.
@@
As for relative-advantage estimation, most existing approaches pool hint-rollouts and naive-rollouts into a single group so that the relative-advantage is dominated by hint-rollouts, with the model learning to directly imitate the off-policy hint distribution rather than exploring with its own policy under hint guidance. Specifically, externally guided hint-rollouts are typically simpler than naive-rollouts, leading to more trajectories with positive advantages that dominate the update signal and bias the policy toward fitting the off-policy distribution. Additionally, hints from external reasoning models are usually much longer than the policy’s own outputs, further exacerbating the imbalance. As shown in Figure 1(b), the average reward of hint-rollouts increases steadily, whereas that of naive-rollouts collapses, causing the model (denoted as Baseline) to eventually lose the ability to perform inference without hints.
Figure 2: Overall comparison between ADHint and baselines. Each axis corresponds to a dataset–metric pair, where p@1 denotes pass@1, and a@8 denotes avg@8. ADHint achieves the best or highly competitive performance on most metrics.
@@
To address these issues, we propose ADHint (ADaptive Hints with Difficulty Priors for Reinforcement Learning), which incorporates difficulty into both the hint-ratio schedule and the relative-advantage estimation to achieve a principled trade-off between exploration and imitation. For the hint-ratio schedule, we propose Adaptive Hint with Sample Difficulty Prior (AH-SDP), which adaptively determines the hint ratio for each sample based on the difficulty prior estimated from the average reward of its naive-rollouts, keeping the resulting hint-rollouts within a moderate difficulty range to provide stable update signals. Furthermore, to prevent large shifts toward the off-policy distribution that trigger entropy collapse, we propose Consistency-based Gradient Modulation (CGM), which downweights the gradients of hint tokens whose entropy deviates significantly from that of the policy-generated continuation. We also introduce Selective Masking for Hint Preservation (Selective Masking), which filters out erroneous update signals from hint-rollouts with negative advantages. For the relative-advantage estimation, we propose Advantage Estimation with Rollout Difficulty Posterior (AE-RDP), which constructs the rollout difficulty posterior from the average rewards of both naive-rollouts and hint-rollouts, and uses it to estimate advantages. Positive naive-rollouts with greater difficulty receive larger advantages, as they provide valuable learning signals that are better aligned with current policy, whereas negative hint-rollouts with lower difficulty are penalized more heavily.
@@
We conduct extensive experiments across both multimodal and text-only settings, multiple model scales, and diverse benchmarks covering mathematical, multi-disciplinary, and logical reasoning. As shown in Figure 2, ADHint consistently yields substantial gains in pass@1 and avg@8, demonstrating synergistic improvements in both reasoning ability and out-of-distribution generalization. Experiments on the recent, stronger backbone Qwen3-VL-8B-Instruct [yang2025qwen3] further validate the effectiveness of ADHint. Our main contributions are summarized as follows:
@@
  • We reveal that difficulty is a crucial signal for both hint-ratio schedule and relative-advantage estimation, and that neglecting it results in unstable learning and overfitting to the off-policy distribution.
@@
  • We propose ADHint, which explicitly exploits sample difficulty priors and rollout difficulty posteriors for hint-ratio schedule and relative-advantage estimation, leading to a better balance between exploration and imitation.
@@
  • Extensive experiments across diverse settings consistently demonstrate ADHint’s strong performance, delivering significant improvements in pass@1 and avg@8 on multi-domain benchmarks.
@@ 2.1 Multimodal Reinforcement Learning
GRPO [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages via intra-group sampling, which has become a cornerstone for recent reasoning models [yang2025qwen3, zeng2025glm, team2025kimi, team2025longcat, ye2025limo, wang2025reinforcement, tan2025towards]. In parallel, the multimodal community has explored various strategies to enhance MLLM reasoning, including refining reasoning styles [wang2025vl, zheng2025deepeyes, su2025pixel], improving rollout procedures for training robustness [liu2025noisyrollout, wang2025perception], redesigning reward functions [liu2025visionreasoner, liu2025visual], and data augmentation [liang2025modomodomultidomaindatamixtures, li2025truth]. Despite these advances, RL algorithms primarily focus on unlocking the latent potential of pretrained models and often fail to expand their fundamental capability boundary.
2.2 Reinforcement Learning with Off-Policy Data
Recent work explores incorporating off-policy data into RL-based post-training. Methods that directly integrate off-policy data into SFT or RL objectives [yan2025learning, ma2025learning, dong2025rl, fu2025srft] often struggle with the substantial distribution mismatch between off-policy data and the on-policy model. Hint-based methods alleviate this by guiding the model’s reasoning with partial hints extracted from off-policy data while preserving its own exploration. UFT [liu2025uft] and Prefix-RFT [huang2025blending] employ temporally annealed hint ratios, while StepHint [zhang2025stephint] mixes hint-rollouts from multiple hint ratios with naive-rollouts and ground-truth into a single group. Other methods reserve hints only for the hardest samples and determine the hint ratio via uniform sampling [chen2025data], binary search [zhang2025bread], or progressive search [Huang_2025_ICCV, liu2025ghpo]. In this work, we propose ADHint, which explicitly incorporates difficulty into both the hint-ratio schedule and the relative-advantage estimation to address the unstable learning and excessive imitation observed in existing hint-based methods. More detailed discussions are provided in Appendix B.
@@ 3 Preliminary
Figure 3: Overview of ADHint. For each sample, ADHint successively infers twice to obtain naive-rollouts and hint-rollouts, where the hint ratio of the latter is scheduled based on the difficulty evaluated from the former. It then leverages the relative difficulty of naive-rollouts and hint-rollouts to estimate their respective advantages, and modulates the gradient of each hint token based on the consistency of entropy.
@@ 3.1 Group Relative Policy Optimization
$\mathcal{J}_{\text{GRPO}}(\pi_{\theta})={}\mathbb{E}_{q\sim Q,{\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}}$
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Biggl\{\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i}\;-\beta\,D_{\mathrm{KL}}\bigl(\pi_{\theta}\,\|\,\pi_{\text{ref}}\bigr)\Biggr\}.$
@@
$\nabla_{\theta}\mathcal{J}^{*}_{\text{GRPO}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}$
$\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\hat{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
@@ 4.1 Overview
As shown in Figure 3, ADHint comprises four crucial modules: (1) Adaptive Hint with Sample Difficulty Prior (AH-SDP), where the policy model successively performs two inferences to obtain naive-rollouts and hint-rollouts, and accordingly schedules the hint ratio for the latter based on the sample difficulty prior evaluated from the former; (2) Advantage Estimation with Rollout Difficulty Posterior (AE-RDP), which uses the rollout difficulty posterior to estimate relative advantages for naive-rollouts and hint-rollouts, covering both correct and wrong outcomes; (3) Consistency-based Gradient Modulation (CGM), where the gradient of each hint token is modulated according to the consistency of its entropy distribution with that of the continuation; and (4) Selective Masking for Hint Preservation (Selective Masking), which masks the update signals of hint tokens in hint-rollouts that remain incorrect even with guidance. Additionally, Algorithm 1 summarizes the overall procedure of ADHint.
@@ 4.2 Adaptive Hint with Sample Difficulty Prior
Unlike existing methods, we use the difficulty prior evaluated from each sample’s naive-rollouts to schedule the hint ratio, thereby keeping hint-rollouts within a moderate-difficulty regime (Figure 5(a)) and providing low-variance update signals at every step. For each query $q$, we define the difficulty score of naive-rollouts as
@@
where $w_{\text{max}}$ and $w_{\text{min}}$ denote the maximum and minimum ratio hyperparameters, and $\sigma\sim\mathcal{U}(-R,R)$ is uniformly distributed noise.
@@
$\bar{H}_{i}\;=\;\frac{1}{\,|o_{i,>h}|\,}\sum_{t=h+1}^{|o_{i}|}\mathcal{H}\!\big(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\big).$ (4)
@@
$k_{i,t}\;=\;\begin{cases}g\!\left(\dfrac{H_{i,t}}{\bar{H}_{i}}\right),&1\leq t\leq h,\\[6.0pt] 1,&t>h,\end{cases}$
@@
$g(x;\alpha)\;=\;\begin{cases}\displaystyle\!\sin\!\Big(\tfrac{\pi}{2}\cdot\tfrac{x-\alpha}{1-\alpha}\Big),&\alpha\leq x\leq 1,\\[8.0pt] \displaystyle\!\cos\!\Big(\tfrac{\pi}{2}\cdot\tfrac{x-1}{(1/\alpha)-1}\Big),&1<x\leq 1/\alpha,\\[8.0pt] 0,&\text{otherwise}.\end{cases}$
@@
$k_{i,t}\;=\;\begin{cases}0,&1\leq t\leq h\ \text{and}\ \hat{A}_{i}\leq 0,\\[4.0pt] g\!\left(\dfrac{H_{i,t}}{\bar{H}_{i}}\right),&1\leq t\leq h\ \text{and}\ \hat{A}_{i}>0,\\[6.0pt] 1,&t>h.\end{cases}$ (5)

### v2

@@ Abstract
To address the limited capability expansion and low sample efficiency of Reinforcement Learning (RL), recent methods have integrated “hints” into post-training, which are prefix segments of complete reasoning trajectories, aiming for powerful knowledge expansion and reasoning generalization. However, existing hint-based RL methods often neglect the role of difficulty in the hint-ratio schedule and relative-advantage estimation, resulting in unstable learning and excessive imitation of off-policy hints. To address this, we propose ADHint, which explicitly integrates difficulty into both processes to achieve a better trade-off between exploration and imitation. Specifically, we propose Adaptive Hint with Sample Difficulty Prior, which evaluates the difficulty of each sample under the current policy to schedule an appropriate hint ratio for rollout generation. Furthermore, we introduce Consistency-based Gradient Modulation alongside Selective Masking for Hint Preservation, which jointly modulate token-level gradients within hints to prevent biased and destructive updates. Additionally, we propose Advantage Estimation with Rollout Difficulty Posterior, which leverages the relative difficulty of rollouts with and without hints to compute their respective advantages, yielding more balanced updates. Extensive experiments across diverse modalities, model scales, model families, and domains demonstrate that ADHint achieves superior reasoning capabilities and out-of-distribution generalization. Code and datasets will be made publicly available upon paper acceptance.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) [yang2024qwen2, yang2025qwen3, li2025system, zeng2025glm] and Multimodal Large Language Models (MLLMs) [bai2025qwen3vltechnicalreport, liu2025visual, huang2025vision, wang2025internvl3] highlight the potential to enhance reasoning capabilities by large-scale Reinforcement Learning with Verifiable Rewards (RLVR) [shao2024deepseekmath, guo2025deepseek, yu2025dapo], such as Proximal Policy Optimization (PPO) [schulman2017proximal] and Group Relative Policy Optimization (GRPO) [shao2024deepseekmath, guo2025deepseek]. However, current on-policy RLVR faces two major challenges: (1) Limited capability expansion: RLVR is inherently bounded by the base model. It primarily amplifies existing behaviors and refines known reasoning chains, rather than instilling genuinely novel reasoning abilities beyond its initial capability boundaries [chusft, li2025system]. (2) Low sample efficiency: The learning process is bottlenecked by the current policy’s performance, yielding critically sparse reward signals that render hard samples difficult to exploit [wu2025templaterlstructuredtemplateguidedreinforcement, xi2025bapostabilizingoffpolicyreinforcement].
@@
To mitigate these limitations, recent methods [Huang_2025_ICCV, liu2025uft] introduce hints, defined as prefix segments of complete reasoning trajectories, which guide the policy to explore the continuation, thereby acquiring novel capabilities and solving challenging problems. Specifically, these methods assign fixed [zhang2025stephint] or time-varying [liu2025uft, huang2025blending] hint ratios to all samples (or only to the hardest samples [Huang_2025_ICCV, liu2025ghpo]), and then utilize hint-rollouts derived from hint-guided reasoning together with naive-rollouts generated without hints for group relative advantage estimation [zhang2025stephint, huang2025blending].
@@
Despite their empirical successes, existing hint-based RL methods are sensitive to sample-difficulty in the hint-ratio schedule and rollout-difficulty in the relative-advantage estimation. Specifically, most previous methods apply a uniform hint ratio to both easy and hard samples for model reasoning, and then generate hint-rollouts with various difficulties. Policy learning from such heterogeneous rollouts could introduce high-variance model optimization, and thus undermine training stability [bengio2009curriculum, parashar2025curriculum, li2025adacurl]. As shown in Figure 1(a), the model with annealing hint ratio (denoted as Baseline) suffers from training collapse since its entropy rises sharply at the end of training, indicating insufficient absorption of external knowledge and a drift in the model’s own policy. In contrast, when incorporating the hint-ratio schedule with sample-difficulty (our method, discussed later), the entropy converges stably, ensuring that the model internalizes robust reasoning abilities rather than superficial patterns within off-policy hints.
@@
As for relative-advantage estimation, most existing approaches pool hint-rollouts and naive-rollouts into a single group so that the relative-advantage is dominated by hint-rollouts, with the model learning to directly imitate the off-policy hint distribution rather than exploring with its own policy under hint guidance. Specifically, externally guided hint-rollouts are typically simpler than naive-rollouts, leading to more trajectories with positive advantages that dominate the update signal and bias the policy toward fitting the off-policy distribution. Additionally, hints from external reasoning models are usually much longer than the policy’s own outputs, further exacerbating this imbalance. As shown in Figure 1(b), the average reward of hint-rollouts increases steadily, whereas that of naive-rollouts collapses, causing the model (denoted as Baseline) to eventually lose the ability to perform inference without hints.
(a) Performance on Qwen2.5-VL-7B.
(b) Gains across scenarios.
Figure 2: Performance Comparison and Generalization. (a) Radar chart on Qwen2.5-VL-7B-Instruct. Each axis represents a dataset-metric pair (p@1: pass@1, a@8: avg@8), where ADHint consistently achieves the best or highly competitive performance. (b) Absolute gains over the best baselines. ADHint demonstrates robust improvements across diverse modalities, model scales, model families, and domains.
@@
To address these issues, we propose ADHint (ADaptive Hints with Difficulty Priors for Reinforcement Learning), which incorporates difficulty into both the hint-ratio schedule and the relative-advantage estimation to achieve a principled trade-off between exploration and imitation. For the hint-ratio schedule, we propose Adaptive Hint with Sample Difficulty Prior (AH-SDP), which adaptively determines the hint ratio for each sample based on a difficulty prior estimated from the average reward of its naive-rollouts, maintaining the resulting hint-rollouts within a moderate difficulty range to provide stable update signals. Furthermore, to prevent drastic shifts toward the off-policy distribution that trigger entropy collapse, we propose Consistency-based Gradient Modulation (CGM), which downweights the gradients of hint tokens whose entropy deviates significantly from that of the policy-generated continuation. We also introduce Selective Masking for Hint Preservation (Selective Masking), which discards erroneous update signals from hint-rollouts with negative advantages. For the relative-advantage estimation, we propose Advantage Estimation with Rollout Difficulty Posterior (AE-RDP), which constructs the rollout difficulty posterior from the average rewards of both naive-rollouts and hint-rollouts, and uses it to estimate advantages. Positive naive-rollouts with greater difficulty receive larger advantages, as they provide valuable learning signals better aligned with the current policy, whereas negative hint-rollouts with lower difficulty are penalized more heavily.
@@
We conduct extensive experiments across multiple modalities, model scales, model families, domains, and tasks. As shown in Figure 2, ADHint consistently yields substantial gains, demonstrating synergistic improvements in both reasoning ability and out-of-distribution generalization. Our main contributions are summarized as follows:
@@
  • We reveal that difficulty is a crucial signal for both the hint-ratio schedule and relative-advantage estimation, and that neglecting it results in unstable learning and overfitting to the off-policy distribution.
@@
  • We propose ADHint, which explicitly exploits sample difficulty priors and rollout difficulty posteriors for the hint-ratio schedule and relative-advantage estimation, leading to a better balance between exploration and imitation.
@@
  • Extensive experiments across diverse settings consistently demonstrate ADHint’s superiority, delivering robust and significant performance gains across comprehensive multi-domain benchmarks.
@@ 2.1 Multimodal Reinforcement Learning
GRPO [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages via intra-group sampling, serving as a cornerstone for recent reasoning models [yang2025qwen3, zeng2025glm, team2025kimi, team2025longcat, ye2025limo, wang2025reinforcement, tan2025towards]. In parallel, the multimodal community has explored various strategies to enhance MLLM reasoning, including refining reasoning styles [wang2025vl, zheng2025deepeyes, su2025pixel], improving rollout procedures for training robustness [liu2025noisyrollout, wang2025perception], redesigning reward functions [liu2025visionreasoner, liu2025visual], and augmenting training data [liang2025modomodomultidomaindatamixtures, li2025truth]. Despite these advances, RL algorithms mainly unlock the latent potential of pretrained models and often fail to expand their fundamental capability boundary.
2.2 Reinforcement Learning with Off-policy Hints
Recent works incorporate off-policy hints into RL-based post-training by providing partial solution prefixes from off-policy data to guide the model’s reasoning while preserving on-policy exploration. UFT [liu2025uft] and Prefix-RFT [huang2025blending] employ temporally annealed hint ratios, whereas StepHint [zhang2025stephint] mixes hint-rollouts at multiple hint ratios with naive-rollouts and ground-truth trajectories within a single group. SEELE [li2025staying] introduces learnable parameters to predict instance-specific hint ratios in a multi-round training scheme. Other methods reserve hints exclusively for the most challenging samples and determine the hint ratio via uniform sampling [chen2025data], binary search [zhang2025bread], or progressive search [Huang_2025_ICCV, liu2025ghpo]. In this work, we propose ADHint, which explicitly incorporates difficulty into both the hint-ratio schedule and the relative-advantage estimation to address the unstable learning and excessive imitation observed in existing hint-based methods. More detailed discussions are provided in Appendix C.
Figure 3: An overview of ADHint. For each sample, ADHint successively infers twice to obtain naive-rollouts and hint-rollouts, where the hint ratio of the latter is scheduled based on the difficulty evaluated from the former. It then leverages the relative difficulty of naive-rollouts and hint-rollouts to estimate their respective advantages, and modulates the gradient of each hint token based on the consistency of entropy.
@@ 3.1 Group Relative Policy Optimization
$\mathcal{J}_{\text{GRPO}}(\pi_{\theta})=\mathbb{E}_{q\sim Q,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Biggl\{\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i}-\beta\,D_{\mathrm{KL}}\bigl(\pi_{\theta}\,\|\,\pi_{\text{ref}}\bigr)\Biggr\}$
@@
$\nabla_{\theta}\mathcal{J}^{*}_{\text{GRPO}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\hat{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
@@ 4.1 Overview
As shown in Figure 3, ADHint comprises four crucial modules: (1) Adaptive Hint with Sample Difficulty Prior (AH-SDP), where the policy model successively performs two inferences to obtain naive-rollouts and hint-rollouts, and accordingly schedules the hint ratio for the latter based on the sample difficulty prior evaluated from the former; (2) Advantage Estimation with Rollout Difficulty Posterior (AE-RDP), which uses the rollout difficulty posterior to estimate relative advantages for naive-rollouts and hint-rollouts, covering both correct and wrong outcomes; (3) Consistency-based Gradient Modulation (CGM), where the gradient of each hint token is modulated according to the consistency of its entropy distribution with that of the continuation; and (4) Selective Masking for Hint Preservation (Selective Masking), which masks the update signals of hint tokens in hint-rollouts that remain incorrect even with guidance. Additionally, Algorithm 1 summarizes the overall procedure.
@@ 4.2 Adaptive Hint with Sample Difficulty Prior
Unlike existing methods, we use the difficulty prior evaluated from each sample’s naive-rollouts to schedule the hint ratio, thereby keeping hint-rollouts within a moderate-difficulty regime (Figure 4(c)) and providing low-variance update signals at every step. For each query $q$, we define the difficulty score of naive-rollouts as
@@
where $w_{\text{max}}$ and $w_{\text{min}}$ denote the maximum and minimum ratios, and $\sigma\sim\mathcal{U}(-R,R)$ is uniformly distributed noise. We fix $w_{\text{min}}=0$ since easy questions require no hints, thus eliminating it as a hyperparameter.
@@
$\bar{H}_{i}=\frac{1}{|o_{i,>h}|}\sum_{t=h+1}^{|o_{i}|}\mathcal{H}(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})).$ (4)
@@
$k_{i,t}=g(H_{i,t}/\bar{H}_{i})\text{ if }1\leq t\leq h,\text{ and }1\text{ otherwise.}$
@@
$g(x;\alpha)=\begin{cases}\sin\bigl(\frac{\pi}{2}\cdot\frac{x-\alpha}{1-\alpha}\bigr),&\alpha\leq x\leq 1,\\[4.0pt] \cos\bigl(\frac{\pi}{2}\cdot\frac{x-1}{1/\alpha-1}\bigr),&1<x\leq 1/\alpha,\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
@@
$k_{i,t}=\begin{cases}0,&1\leq t\leq h\text{ and }\hat{A}_{i}\leq 0,\\ g(H_{i,t}/\bar{H}_{i}),&1\leq t\leq h\text{ and }\hat{A}_{i}>0,\\ 1,&t>h.\end{cases}$ (5)